# What Makes a Scene ? Scene Graph-based Evaluation and Feedback for Controllable Generation

### v1

@@ Abstract
While text-to-image generation has been extensively studied, generating images from scene graphs remains relatively underexplored, primarily due to challenges in accurately modeling spatial relationships and object interactions. To fill this gap, we introduce Scene-Bench, a comprehensive benchmark designed to evaluate and enhance the factual consistency in generating natural scenes. Scene-Bench comprises MegaSG, a large-scale dataset of one million images annotated with scene graphs, facilitating the training and fair comparison of models across diverse and complex scenes. Additionally, we propose SGScore, a novel evaluation metric that leverages chain-of-thought reasoning capabilities of multimodal large language models (LLMs) to assess both object presence and relationship accuracy, offering a more effective measure of factual consistency than traditional metrics like FID and CLIPScore. Building upon this evaluation framework, we develop a scene graph feedback pipeline that iteratively refines generated images by identifying and correcting discrepancies between the scene graph and the image. Extensive experiments demonstrate that Scene-Bench provides a more comprehensive and effective evaluation framework compared to existing benchmarks, particularly for complex scene generation. Furthermore, our feedback strategy significantly enhances the factual consistency of image generation models, advancing the field of controllable image generation.
@@ 1 Introduction
To evaluate the factual consistency, we employ a structured representation known as a Scene Graph, which has been demonstrated to outperform pure text in image retrieval [22, 27]. A scene graph encodes objects as nodes and their relationships as edges. For textual conditions, scene graphs can be parsed using natural language processing tools such as Scene Parser [37] or large language models (LLMs). For image conditions, scene graphs are generated via Scene Graph Generation (SGG) models [52, 58, 48, 10] or multimodal LLMs. Leveraging this structured representation, we introduce a novel evaluation metric, SGScore, which quantifies the factual consistency between generated images and their corresponding scene graphs. SGScore evaluates Object Recall by verifying the presence of nodes and Relation Recall by assessing the accuracy of edges within the scene graph. To adapt different domains and handle the extensive vocabulary inherent in generated images, we utilize a multimodal LLM to perform these evaluations instead of relying on a pre-trained SGG model to convert images into scene graphs. Thanks to the chain-of-thought reasoning and zero-shot capabilities of the multimodal LLM, SGScore can effectively distinguish between images depicting subtle differences, as shown in Fig. 1.
When applying the new metric to benchmark different generative models, a critical bottleneck is the lack of large-scale datasets annotated with scene graphs, which are essential for fair and comprehensive comparisons. Existing datasets, such as Visual Genome (VG) [27] and COCO-Stuff [5], are relatively small (e.g., only 5k and 2k images for testing, respectively), and their inherent long-tail distributions lead to unfair and biased evaluations. As a result, we develop MegaSG, a large-scale dataset comprising one million images richly annotated with scene graphs that capture a wide range of objects and their complex relationships. MegaSG enables models to be trained and evaluated on diverse scenarios, from simple to highly intricate scenes, thus overcoming the limitations of previous datasets that were constrained by small-scale and biased distributions.
@@
  • We introduce Scene-Bench, a comprehensive and large-scale benchmark for evaluating factual consistency using scene graphs. Scene-Bench includes MegaSG, a dataset with one million images annotated with scene graphs, and a novel evaluation metric, SGScore, which explicitly measures factual consistency by assessing the accuracy of object presence and relationships in generated images.
@@ 2 Related Work
Text-to-Image Generation. The field of text-to-image generation has seen significant advancements with the transition from GANs [17] to diffusion models. Early GAN-based methods like StackGAN [59] and AttnGAN [53] generated images from textual descriptions but often struggled with image quality and diversity. The introduction of diffusion models marked a substantial improvement. Models such as DALL-E [42], GLIDE [38], and Stable Diffusion [44] have achieved high-quality image synthesis with better text-image alignment by iteratively refining images from noise, conditioned on text prompts. Despite their success, these models face challenges in generating complex scenes involving multiple objects and ensuring consistency in object relationships.
@@
Scene Graph-to-Image Generation. Scene graphs offer a structured representation of objects and their relationships, providing a promising scheme for controllable image generation. Johnson et al. [23] introduced SG2Im, a model that generates images from scene graphs using graph convolutional networks (GCN) and conditional GANs. Ashual and Wolf [3] extended this approach by incorporating more detailed scene representations and object attributes. Recent methods have integrated scene graphs with diffusion models to enhance compositionality [35, 33]. However, these approaches often require large-scale scene graph datasets and rely on additional guidance like bounding boxes [13] or specialized graph encoders, which cannot be adapted to open vocabulary. The evaluation of these models remains a challenge due to the lack of metrics that capture the fidelity of object relationships in generated images.
Large Language Models in Image Generation. The integration of LLMs has opened new possibilities in image generation. Recent works have leveraged LLMs to enhance compositionality and controllability in image synthesis. For example, LayoutGPT [15] utilizes LLMs as visual planners to generate layouts from textual descriptions, improving user controllability. Similarly, methods like RPG [55] and Complex Diffusion [34] leverage the chain-of-thought reasoning capabilities of LLMs to decompose complex prompts into simpler tasks, aiding in the generation of complex scenes with multiple objects and relationships. However, their potential for providing feedback to iteratively refine scene graph-based generation has not been fully explored.
@@
Dataset Images Vocabulary Diversity Test
Samples Triplets Complexity Balanced
VG [27] 108k 179 objects, 49 relations 5,096 12.8k
COCO-Stuff [5] 4.5k 171 objects, 6 relations 2,048 22.7k
MegaSG 1M 775 objects, 122 relations 50,000 275k
Table 1: Comparison of MegaSG with existing scene graph datasets.
@@ 3 Scene-Bench
Scene-Bench is designed for evaluating and enhancing the factual consistency in generating natural scenes, in terms of objects and relationships. Specifically, Scene-Bench consists of a large-scale dataset of scene graphs, and an autonomous evaluation pipeline.
@@ 3.1 MegaSG: a large-scale dataset of scene graphs
Creation of the Dataset. Due to the complexity and high cost of manual annotation, existing scene graph datasets, such as Visual Genome [27], are relatively small in scale (e.g., only 5k images are prepared for the test set [23]). The limited size and inherent long-tail distribution make these datasets inadequate for studying diffusion models across diverse scene scenarios. To address this limitation and build a large-scale scene graph dataset, we leverage the chain-of-thought reasoning capabilities of multimodal large language models (LLMs) in combination with pre-existing object detection datasets. Specifically, we collect 1 million images from COCO [8], Object365 [46], and Open Images v6 [28], which offer rich object categories and bounding boxes. These datasets are ideal for generating large-scale scene graphs efficiently.
As shown in Fig. 2, the multimodal LLM takes an image with object categories and bounding boxes as input and generates a scene graph that describes the relationships between objects. For instance, in the provided example, the model identifies a “person” kicking a “sports ball” and another “person” nearby, resulting in the following scene graph: {“source”: “person.2”, “target”: “sports ball.1”, “relation”: “kicking”}, {“source”: “person.2”, “target”: “person.3”, “relation”: “near”}. This structured representation captures both the objects and their spatial and relational interactions.
Scene Diversity. To better understand the behavior of diffusion models in diverse scene scenarios, we utilized an LLM (e.g., Gemini 1.5 Flash [43]) to classify the MegaSG dataset into two main themes: People-Centric and Non-People-Centric. The People-Centric theme includes fine-grained categories such as Social Interaction, Individual Activities, Work / Occupation, Travel / Exploration, Sports & Recreation, Performance / Entertainment, Daily Life. For the Non-people-Centric theme, we identified categories like Nature, Urban / Built, Objects, and Abstract / Artistic. The hierarchical distribution of these categories is illustrated in Fig. 2 (b). And samples are shown in the figure (see Supplementary materials), showcasing the diversity and range of scenarios covered in the dataset.
@@
To quantitatively verify the quality of MegaSG, we evaluate Scene Graph Generation (SGG) models trained on different datasets. For instance, the OvSGTR (Swin-B) [10] model trained on MegaSG achieves a zero-shot performance recall of 45.71% (R@50, PredCls mode) on the VG150 test set, outperforming models trained on smaller datasets like COCO Caption data (see Supplementary Materials). This improvement reflects the high-quality scene graph annotations of MegaSG, enabling further exploration of training SGG models on MegaSG or generating images from scene graphs on MegaSG.
@@ 4 Scene Graph Feedback
Specifically, given a scene graph $G=(V,E)$, we first perform scene composition using an LLM, in which nodes and edges are seamlessly integrated into a $prompt$ for an exact scene. This $prompt$ results in an initial image $I_{0}$ through the diffusion model $f_{D}$. With the input scene graph $G$ and the generated image $I_{0}$, a multimodal LLM $f_{M}$ has been applied to evaluate the presence of objects and relationships. The missing objects and relationships are constructed as a missing graph $G_{\text{miss}}$. If discrepancies exists, e.g., $G_{\text{miss}}\neq(\emptyset,\emptyset)$, we will generate a reference image $I_{1}$ conditioned on the $G_{\text{miss}}$ as does in generating $I_{0}$ conditioned on $G$. To generate the final output image, we use IP-Adapter [56] to integrate $prompt$, $I_{0}$, and $I_{1}$, in which the cross attention process can be formulated as
@@ 5.1 Experimental Setup
Datasets. We use the Visual Genome dataset [27] following the data splits from SG2Im [23], and the proposed MegaSG dataset. For a fair comparison, we balance samples for testing based on Scene Diversity and Scene Complexity (see supplementary materials).
Metrics. We employ common metrics such as Inception Score (IS) [45], Fréchet Inception Distance (FID) [19], and CLIPScore [18]. Additionally, we introduce ObjectRecall, RelationRecall, and SGScore (see Sec. 3.2 and supplementary materials for details).
Scene Graph Representation and Complexity. For text-to-image models, we format the scene graphs as “{subject} {predicate} {object}” (e.g., cat sitting on desk, dog near chair). Based on Scene complexity defined as Eq. 6, we categorize the scene complexity into three levels: simple, medium, and hard (details in supplementary materials).
LLM. We use Gemini 1.5 Flash [43] (cutoff November 2024) as the multimodal LLM in our experiments. We also report results using local multimodal LLMs like LLaVA [32] in supplementary materials.
@@ 5.2 Evaluation of Scene-Bench
Impact of Scene Complexity. Beyond the three coarse levels of scene complexity, we analyze the effect of scene complexity ranging from 1 to 10 on model performance. For more details, please see Supplementary materials.
Refer to captionRefer to captionRefer to captionRefer to caption62.5Refer to caption50.0Refer to caption50.0Refer to caption62.5Refer to caption100.0ComposableSD v1.5SD v2.1SDXLOursScene GraphRefer to caption50.0Refer to caption50.0Refer to caption50.0Refer to caption50.0Refer to caption100.0Refer to caption19.6Refer to caption66.0Refer to caption73.2Refer to caption59.8Refer to caption100.0
@@ 5.5 Human Evaluation
We perform a human evaluation to demonstrate how SGScore enhances factual consistency verification. To this end, annotators are tasked with answering 1,000 four-to-one questions (examples can be found in Supplementary materials). Each question presents an original image alongside four generated images from different models, and annotators are asked to choose the image that best aligns with the original in terms of object presence and relationships. The results, as shown in Fig. 5, demonstrate a clear preference for our model over others, both by human annotators and machine selections. This preference indicates a higher factual consistency of images generated by our model, as determined by SGScore.
@@ 6 Conclusion
In this work, we introduced Scene-Bench, a comprehensive and large-scale benchmark designed to evaluate factual consistency in the generation of natural scenes. Scene-Bench includes the large-scale MegaSG dataset and a novel evaluation metric, SGScore, which assesses object presence and relationship accuracy using multimodal large language models. The proposed scene graph feedback leverages explicit evaluations to iteratively refine generated images, significantly improving factual consistency. Extensive experiments demonstrate that Scene-Bench provides a rigorous evaluation framework, particularly excelling in complex scene scenarios where traditional metrics fall short. By effectively addressing the challenges of accurately modeling spatial relationships and object interactions, the proposed scheme advances the field of controllable generation and sets a new benchmark for future research.

### v2

@@ Abstract
While text-to-image generation has been widely explored, synthesizing images from scene graphs remains relatively underexplored due to challenges in capturing complex spatial relationships and object interactions. To bridge this gap, we introduce Scene-Bench, a comprehensive benchmark for evaluating and enhancing factual consistency in natural scene generation. Scene-Bench comprises MegaSG, a large-scale dataset of one million images annotated with detailed scene graphs, enabling extensive training and fair comparisons across diverse and intricate scenes. In addition, we propose SGScore, a novel evaluation metric that leverages the reasoning capabilities of multimodal large language models to assess both object presence and relationship accuracy, thereby providing a more precise measure of factual consistency than traditional metrics such as FID and CLIPScore. Furthermore, our scene graph feedback pipeline iteratively refines generated images by identifying and correcting discrepancies between the intended scene graph and the output. Extensive experiments demonstrate that Scene-Bench offers an effective evaluation framework for complex scene generation, and our feedback strategy significantly improves the factual consistency of image generation models, advancing the field of controllable image generation.
@@ 1 Introduction
To evaluate the factual consistency, we employ a structured representation known as a Scene Graph, which has been demonstrated to outperform pure text in image retrieval [26, 31]. A scene graph encodes objects as nodes and their relationships as edges. For textual conditions, scene graphs can be parsed using natural language processing tools such as Scene Parser [44] or large language models (LLMs). For image conditions, scene graphs are generated via Scene Graph Generation (SGG) models [64, 70, 58, 11] or multimodal LLMs. Leveraging this structured representation, we introduce a novel evaluation metric, SGScore, which quantifies the factual consistency between generated images and their corresponding scene graphs. SGScore evaluates Object Recall by verifying the presence of nodes and Relation Recall by assessing the accuracy of edges within the scene graph. To adapt different domains and handle the extensive vocabulary inherent in generated images, we utilize a multimodal LLM to perform these evaluations instead of relying on a pre-trained SGG model to convert images into scene graphs. Thanks to the reasoning and zero-shot capabilities of the multimodal LLM, SGScore can effectively distinguish between images depicting subtle differences, as shown in Fig. 1.
When applying the new metric to benchmark different generative models, a critical bottleneck is the lack of large-scale datasets annotated with scene graphs, which are essential for fair and comprehensive comparisons. Existing datasets, such as Visual Genome (VG) [31] and COCO-Stuff [6], are relatively small (e.g., only 5k and 2k images for testing, respectively), and their inherent long-tail distributions lead to biased evaluations. As a result, we develop MegaSG, a large-scale dataset comprising one million images richly annotated with scene graphs that capture a wide range of objects and their complex relationships. MegaSG enables models to be trained and evaluated on diverse scenarios, from simple to highly intricate scenes, thus overcoming the limitations of previous datasets that were constrained by small-scale and biased distributions.
@@
  • We introduce Scene-Bench, a comprehensive and large-scale benchmark for evaluating factual consistency in scene graph-to-image generation. Scene-Bench includes MegaSG, a dataset with one million images annotated with scene graphs, and a novel evaluation metric, SGScore, which explicitly measures factual consistency by assessing the accuracy of object presence and relationships in generated images.
@@ 2 Related Work
Text-to-Image Generation (T2I). The field of text-to-image generation has seen significant advancements with the transition from GANs [20] to diffusion models. Early GAN-based methods like StackGAN [71] and AttnGAN [65] generated images from textual descriptions but often struggled with image quality and diversity. The introduction of diffusion models marked a substantial improvement. Models such as DALL-E [50], GLIDE [45], and Stable Diffusion [52] have achieved high-quality image synthesis with better text-image alignment by iteratively refining images from noise, conditioned on text prompts. Despite their success, these models face challenges in generating complex scenes involving multiple objects and ensuring consistency in object relationships.
@@
Scene Graph-to-Image Generation (SG2IM). Scene graphs offer a structured representation of objects and their relationships, providing a promising scheme for controllable image generation. Johnson et al. [27] introduced SG2Im, a model that generates images from scene graphs using graph convolutional networks (GCN) and conditional GANs. Ashual and Wolf [3] extended this approach by incorporating more detailed scene representations and object attributes. Recent methods have integrated scene graphs with diffusion models to enhance compositionality [42, 16, 56, 40, 62]. However, these approaches often require training on large-scale scene graph datasets and rely on additional guidance like bounding boxes(e.g., [16]) or specialized graph encoders (e.g., [66]), which cannot be adapted to open-vocabulary scenarios. More importantly, evaluating these models remains challenging due to the lack of metrics that effectively capture scene-level fidelity, including object presence and relationship accuracy in generated images.
LLMs in Image Generation. The integration of LLMs has opened new possibilities in image generation. Recent works [18, 63, 67, 37] have leveraged LLMs to enhance compositionality and controllability in image synthesis. For example, LayoutGPT [18] utilizes LLMs as visual planners to generate layouts from textual descriptions, improving user controllability. Similarly, methods like RPG [67] and Complex Diffusion [41] leverage the reasoning capabilities of LLMs to decompose complex prompts into simpler tasks, aiding in the generation of complex scenes with multiple objects and relationships. However, their potential for providing feedback to iteratively refine scene graph-based generation has not been fully explored.
@@
Dataset Images Obj./Rel. Test
Samples Triplets Balanced
VG [31] 108k 179 / 49 5,096 12.8k
COCO-Stuff [6] 4.5k 171 / 6 2,048 22.7k
MegaSG 1M 775 / 122 50,000 275k
Table 1: Comparison of MegaSG with widely-used scene graph datasets. “Obj./Rel.” denotes the number of object and relationship categories, while “Balanced” indicates whether the dataset is balanced with respect to scene complexity. The VG statistics follow the pioneering work [27] in SG2IM, and almost all subsequent works adopt this setting.
@@ 3 Scene-Bench
Scene-Bench is a comprehensive benchmark that evaluates and enhances the factual consistency of natural scene generation from scene graphs by rigorously verifying both object presence and inter-object relationships. Specifically, Scene-Bench consists of a large-scale dataset of scene graphs, and an autonomous evaluation pipeline. The overview of Scene-Bench is shown in Fig. 2.
@@ 3.1 MegaSG: a large-scale dataset of scene graphs
Creation of the Dataset. Due to the complexity and high cost of manual annotation, existing scene graph datasets, such as Visual Genome [31], are relatively small in scale (e.g., only 5k images are prepared for the test set [27]). The limited size and inherent long-tail distribution make these datasets inadequate for studying diffusion models across diverse scene scenarios. To address this limitation and build a large-scale scene graph dataset, we leverage the reasoning capabilities of multimodal large language models (LLMs) in combination with pre-existing object detection datasets. Specifically, we collect 1 million images from COCO [9], Object365 [55], and Open Images v6 [32], which offer rich object categories and bounding boxes. These datasets are ideal for generating large-scale scene graphs efficiently. For additional details, please see Appendix A of the SMs  [^1] Throught this paper, SMs refers to supplementary materials..
Dataset Quality. To quantitatively verify the quality of MegaSG, we evaluate state-of-the-art Scene Graph Generation (SGG) models trained on different datasets. For instance, the OvSGTR (Swin-B) [11] model trained on MegaSG achieves a zero-shot performance recall of 45.71% (R@50, PredCls mode) on the VG150 test set, outperforming models trained on smaller datasets like COCO Caption data (see Tab. 5 of SMs). This improvement reflects the high-quality scene graph annotations of MegaSG, enabling further exploration of training SGG models on MegaSG or generating images from scene graphs on MegaSG.
Scene Diversity. To better understand the behavior of diffusion models in diverse scene scenarios, we utilized an LLM (e.g., Gemini 1.5 Flash [51]) to classify the MegaSG dataset into two main themes: People-Centric and Non-People-Centric. The People-Centric theme includes fine-grained categories such as Social Interaction, Individual Activities, Work / Occupation, Travel / Exploration, Sports & Recreation, Performance / Entertainment, Daily Life. For the Non-people-Centric theme, we identified categories like Nature, Urban / Built, Objects, and Abstract / Artistic. The hierarchical distribution of these categories is illustrated in Fig. 2 (b). And samples are shown in the figure (see Fig. 5 of SMs), showcasing the diversity and range of scenarios covered in the dataset.
@@ 4 Scene Graph Feedback
Specifically, given a scene graph $G=(V,E)$, we first perform scene composition using an LLM (see Tab. 6 in the SMs), in which nodes and edges are seamlessly integrated into a $prompt$ for an exact scene. This $prompt$ results in an initial image $I_{0}$ through the diffusion model $f_{D}$. With the input scene graph $G$ and the generated image $I_{0}$, a multimodal LLM $f_{M}$ has been applied to evaluate the presence of objects and relationships. The missing objects and relationships are constructed as a missing graph $G_{\text{miss}}$. If discrepancies exists, e.g., $G_{\text{miss}}\neq(\emptyset,\emptyset)$, we will generate a reference image $I_{1}$ conditioned on the $G_{\text{miss}}$ as does in generating $I_{0}$ conditioned on $G$. To generate the final output image, we use IP-Adapter [68] to integrate $prompt$, $I_{0}$, and $I_{1}$, in which the cross attention process can be formulated as
@@ 5.1 Experimental Setup
Datasets. We use the Visual Genome dataset [31] following the data splits from SG2Im [27], and the proposed MegaSG dataset. For a fair comparison, we balance samples for testing based on Scene Diversity and Scene Complexity (see SMs).
Metrics. We employ common metrics such as Inception Score (IS) [53], Fréchet Inception Distance (FID) [22], and CLIPScore [21]. Additionally, we introduce ObjectRecall, RelationRecall, and SGScore (see Sec. 3.2 and Sec. B.1 of the SMs for details).
Scene Graph Representation. For text-to-image (T2I) models that condition on a sentence, we encode scene graphs in the format “{subject} {predicate} {object}” (e.g., cat sitting on desk, dog near chair). For the prompt that converts a scene graph into a consistent description (i.e., scene composition in Sec. 4), please refer to Sec. B.1 of the SMs.
Scene Complexity. Based on Scene complexity defined as Eq. 1, we categorize the scene complexity into three levels: simple, medium, and hard (details in Sec. B.1 of the SMs).
LLM. We use Gemini 1.5 Flash [51] (cutoff November 2024) as the multimodal LLM in our experiments. We also report results using local multimodal LLMs like LLaVA [39] in Sec. C.1 of the SMs.
@@ 5.2 Evaluation of Scene-Bench
Impact of Scene Complexity. Beyond the three coarse levels, we evaluate model performance across a complexity range from 1 to 10. Detailed experimental results and analysis are provided in Section B.2 of the SMs. Our findings reveal that, although image quality (measured by FID) remains stable, increasing scene complexity significantly degrades the factual consistency of scene representations (measured by SGScore). Notably, our model consistently outperforms competitors by maintaining higher object and relationship recall across all complexity levels.
Refer to captionRefer to captionRefer to caption62.5Refer to caption50.0Refer to caption50.0Refer to caption62.5Refer to caption100.0ComposableSD v1.5SD v2.1SDXLOursScene GraphRefer to caption19.6Refer to caption66.0Refer to caption73.2Refer to caption59.8Refer to caption100.0
@@ 5.3 Evaluation of Scene Graph Feedback
Ablation Study of IP-Adapter. To assess the impact of the additional parameters introduced by the IP-Adapter, we conduct an ablation study. Detailed results are provided in Sec. B.3 of the SMs.
@@ 5.5 Human Evaluation
To validate the efficacy of SGScore in improving the verification of factual consistency and assess the impact of the proposed feedback pipeline, we conducted a human evaluation. Annotators were presented with 1,000 four-to-one comparative queries (see examples in Sec. B.4 of the SMs). Each query displayed an original image alongside four generated images produced by different models. Annotators were instructed to select the image that most accurately preserved the object presence and relationships of the original. As illustrated in Fig. 9 of the SMs, both human judgments and machine-based selections consistently favored our model, thereby confirming its superior factual consistency as measured by SGScore.
6 Discussion
In the field of SG2IM, most existing approaches [27, 3, 66, 16, 56, 40, 62] focus primarily on network architecture design, train on widely used datasets (e.g., Visual Genome and COCO-stuff), and report conventional metrics such as IS, FID, and CLIPScore. However, assessing factual consistency and leveraging identified inconsistency as feedback remain underexplored. To address this gap, we introduce a large-scale scene graph dataset, a novel metric SGScore, and an automatic evaluation pipeline to systematically measure the factual consistency in SG2IM. Based on the evaluation pipeline, we incorporate a training-free feedback pipeline to enhance factual consistency.
One potential concern for the newly introduced dataset, MegaSG, is its lack of attribute annotations (e.g., color, shape, texture) that are present in the VG dataset. To assess the impact of this omission, we present an experimental analysis in Sec. D.2 of the SMs. The results indicate that explicit attribute binding yields a marginal difference. Considering the trade-off between additional annotation costs and minimal gains, we opted to omit attributes in the construction of MegaSG. Beyond this concern, we compare Scene-Bench with existing T2I benchmarks like TIFA [24] and DSG [12] in Sec. D.3 to validate the reliability of SGScore.
@@ 7 Conclusion
In this work, we introduce Scene-Bench, a comprehensive benchmark for evaluating the factual consistency of generating natural scenes from scene graphs. Our benchmark incorporates a large-scale dataset, MegaSG, with a novel metric, SGScore, which quantitatively assesses both the presence of objects and the accuracy of relationships through the reasoning capabilities of multimodal LLMs. Building upon this evaluation, our scene graph feedback mechanism iteratively refines generated images by correcting inconsistencies between the scene graph and the output. This process significantly improves the factual consistency. Extensive experiments demonstrate that Scene-Bench offers a rigorous evaluation framework, especially in complex scenes where traditional metrics fall short. We believe that Scene-Bench will establish a new standard and inspire future research in high-fidelity, controllable generation.
Supplementary Material
Figure 5: Illustration of scene categories in the MegaSG dataset. The image shows various themes, such as People-Centric (e.g., social interaction, individual activities) and Non-People-Centric (e.g., nature, urban environments). The caption is provided for illustrative purposes and generated using BLIP-2 [35], and the scene graph is constructed as described in Sec. 3.1 and Appendix A.
(a) VG Objects
(b) MegaSG Objects
(c) VG Relationships
(d) MegaSG Relationships
Figure 6: Word clouds of objects and relationships in the Visual Genome (VG) and MegaSG datasets. (a) and (b) illustrate the diversity of objects, while (c) and (d) highlight the relationships. The comparison demonstrates MegaSG’s broader vocabulary and richer representation of object-relationship semantics.